# FUSION: Forecast-Embedded Agent Scheduling with Service Incentive Optimization over Distributed Air-Ground Edge Networks

### v1

@@ Abstract
We investigate a forecasting-driven, incentive-compatible service provisioning framework for distributed air–ground integrated edge networks under human–machine coexistence. We consider hybrid players where the computing capacity of edge servers (ESs) are augmented by vehicular–UAV agent pairs (AgPs) that can be proactively dispatched to overloaded hotspots. Unique key challenges should be addressed: highly uncertain spatio–temporal workloads, spatio–temporal coupling between road traffic and UAV capacity, forecast-driven contracting risks, and heterogeneous quality-of-service (QoS) requirements of human users (HuUs) and machine users (MaUs). To cope with these issues, we propose FUSION, a two-stage framework that tightly couples service demand prediction, agent deployment, and task scheduling. In the offline stage, a Pro-LNN module performs intelligent multi-step spatio–temporal demand forecasting at distributed ESs, whose outputs are exploited by an enhanced ant colony optimization–based routing scheme (eACO-VRP) and an auction-based incentive-compatible contracting mechanism (Off-AIC2), to jointly determine ES–AgP contracts and pre-planned service routes. In the online stage, we formulate congestion-aware task scheduling as an potential game among HuUs, MaUs, and heterogeneous ES/UAV providers, and devise a potential-guided best-response dynamics (PG-BR) algorithm that provably converges to a pure-strategy Nash equilibrium. Extensive experiments on both synthetic and real-world traces show that FUSION significantly improves social welfare, latency, resource utilization, and robustness compared with state-of-the-art baselines.
@@
Distributed edge computing, auction, liquid neural network, potential games, heterogeneous users.
@@ I Introduction
With the rapid advancement of edge intelligence and unmanned aerial vehicle (UAV) technologies, a new generation of intelligent applications, such as real-time video analytics, intelligent transportation systems, and large-scale mobile crowdsensing, has flourished across smart cities [1, 2]. These applications are typically computation-intensive (e.g., calling for analyzing large amount of data, such as mode training and inference) and highly delay-sensitive, which imposes stringent requirements on the elasticity and quality of edge services. Although edge servers (ESs) provide proximity computing capabilities, their geographically distributed deployment and limited local resources make them vulnerable to resource shortages and task congestion during peak periods, resulting in severe service latency and performance degradation[3].
To mitigate capacity limitations near the network edge, low-altitude-assisted edge computing has attracted extensive attention [4, 5]. Unlike static ESs, UAVs possess high flexibility and can be rapidly dispatched to hotspot regions to provide additional computing/caching capabilities. However, most existing studies implicitly assume that UAVs can “fly directly” to target regions, while overlooking practical constraints such as limited endurance and high deployment cost[8, 7, 9, 6]. In large-scale urban scenarios with widely distributed ESs and hotspot regions (exhibit intricate spatio-temporal patterns), relying solely on UAV self-flight is insufficient to cover multiple overloaded ESs in time.
To this end, we advocate a air-ground cooperative edge computing paradigm: during periods of intensive demand, agents for service enhancement such as ground vehicles can be used to transport multiple UAVs (vehicles and UAVs are collectively known as “agent pairs (AgPs)” in our paper, each consisting of a vehicle as a carrier and several on-board UAVs) to the vicinities of overloaded ESs. UAVs are then launched from strategically selected locations to deliver temporary computational augmentation for nearby ESs[10, 11]. This not only expands on-demand computing capacity during peak intervals and alleviates ES congestion, but also enables more agile and fine-grained service provisioning for latency-sensitive applications. Despite these benefits, however, the introduction of UAV operations significantly increases system complexity across temporal, spatial, service, and economic dimensions. This complexity gives rise to several fundamental challenges that must be addressed to fully leverage the potential of this hybrid deployment paradigm.
(A) High spatio-temporal demand uncertainty over complex environments with heterogeneous users. Edge service demand exhibits pronounced spatio-temporal variability, and the coexistence of human users (HuUs) and machine users (MaUs) [^1] In this paper, we offer an unique perspective to distinguish between HuUs and MaUs: HuUs refer to individuals whose task performance directly affects their emotion or real-time experience (e.g., drivers using intelligent navigation, players engaged in mobile gaming and AR/VR). In contrast, MaUs denote non-human entities, such as mapping vehicles, delivery robots, or sensing units, whose task performance does not immediately or emotionally influence humans. Consequently, HuUs typically exhibit strict instantaneous latency sensitivity, while MaUs only require task completion within a pre-defined deadline window. These differences lead to fundamentally heterogeneous QoS requirements, and thus utility modelings in subsequent sections. further exacerbates the unpredictability of future workloads. This volatility makes it difficult to determine when, where, and how UAV resources should be dispatched. Static provisioning or purely reactive scheduling often leads to mismatches between supply and demand, manifesting as severe congestion in emerging hotspots and resource underutilization in low-demand regions[13, 12]. The core challenge, therefore, lies in developing accurate and actionable multi-step spatio-temporal demand forecasts based on historical requests and contextual information. Such forecasts are essential for enabling proactive, rather than reactive, UAV deployment decisions; yet achieving reliable predictions under highly dynamic and heterogeneous urban environments remains fundamentally difficult.
(B) Spatio-temporal coupling in complexity of road traffic networks and capacity of UAVs. During peak periods, AgPs can serve overloaded ESs that are geographically distributed. However, vehicular mobility is constrained by road network topology, traffic dynamics, and strict time budgets, while the number of UAVs that each vehicle can deploy is tightly coupled with its visiting sequence and dwell duration at each ES. These spatio-temporal dependencies create a highly entangled decision space. As a result, the underlying vehicle-UAV joint deployment optimization becomes strongly NP-hard, where plans must simultaneously incorporate multi-modal travel costs, strict service time windows, UAV dispatching logistics, and the spatially distributed service rewards linked to relieving ES overload. This tight coupling yields a highly interdependent decision space.
(C) Predictive contracting and incentive-compatible trading. To facilitate timely services, ESs can buy services from AgPs, calling for establishing advance contracts, committing to service capacity and pricing for future peak intervals before those intervals materialize. However, inevitable forecasting errors and dynamically evolving supply-demand conditions introduce substantial risks for both parties: ESs may face unmet demand if the contracted capacity falls short, whereas AgPs may incur stranded capacity or loss from misallocated resources. The central challenge lies in designing an offline trading mechanism that remains individually rational, incentive compatible, and budget feasible[14], while explicitly integrating predictive uncertainty into the contracting process. Achieving such robustness without sacrificing economic efficiency or strategic truthfulness is nontrivial, making it a fundamental difficulty in forecast-driven agents collaboration.
(D) Timely and fairness-aware online coordination under heterogeneous human–machine coexistence. When actual computing demands arrive, HuUs and MaUs may compete for the same pool of resources provided by local ESs and leased UAVs. However, their quality-of-service (QoS) requirements differ fundamentally: HuUs are highly sensitive to instantaneous delay (their emotion can directly be affected), whereas MaUs primarily value timely task completion before deadlines. In proper coordination can lead to individually rational behaviors and unsatisfying services, such as inequitable resource distribution across heterogeneous user groups, recurrent congestion oscillations among ESs, and sub-optimal utilization of UAV-enhanced service capacity. The key challenge is thus to design fairness-aware distributed scheduling with provable convergence guarantees, ensuring both equitable treatment across heterogeneous user types and high system-wide efficiency in human-machine coexistence environments.
@@
These above challenges collectively underscore the core motivation of our work. In this paper, we develop a forecasting-driven agent scheduling (i.e., AgPs formed by vehicles and UAVs) and incentive optimization (FUSION) framework for edge service provision, with two stages involved: seamlessly integrating demand prediction, agent and task scheduling, and incentive design. First, for challenge (A), we design a proactive spatio-temporal demand prediction via liquid neural networks (Pro-LNN), transforms historical service traces and contextual features into multi-step workload forecasts, thereby offering actionable decision support for effective AgP deployment. Then, for challenge (B), we propose an enhanced ant colony optimization for vehicular rout planning (eACO-VRP) based on the predicted demand, jointly models ESs’ demand time windows, vehicular travel cost, and expected deployment rewards, generating cost-effective and executable plans. Third, for challenge (C), we introduce offline auction-based incentive-compatible contracting (Off-AIC2) that matches ESs with proper AgPs, determines payments and rewards, thereby achieving, to a certain extent, individual rationality, incentive compatibility, and budget feasibility. Finally, for challenge (D), we formulate an efficient online potential game among HuUs, MaUs, and heterogeneous ES/UAV, and propose the potential game-based best-response dynamics (PG-BRD) method, which integrates potential-guided dual decomposition with asynchronous best-response dynamics to drive the system towards a pure-strategy Nash equilibrium (NE) under congestion and fairness constraints. To sum up, our principal contributions are summarized as follows:
  • We investigate a distributed air-ground integrated edge network architecture that holistically models the deep coupling among AgPs, and ESs, together with the heterogeneous computing demands arising in human-machine coexistence. To enable seamless, intelligent, and latency-sensitive service delivery, we cast the goal as a unified optimization with NP-hardness, due to intertwined mobility patterns, stochastic workloads, and strategic agent behaviors. To cope with this, we offer an unique view, spanning spatio-temporary demand prediction, agent scheduling, and incentive-compatible coordination, by designing two complementary stages.
  • We first construct a prediction-driven offline optimization stage, where Pro-LNN is designed to perform multi-step spatio-temporal workload prediction by training LNNs; while eACO-VRP exploits the prediction to jointly consider ES time windows, travel cost, and expected rewards, producing cost-effective and practically executable vehicular routings as well as UAV pre-deployment plans. Besides, during this stage, we investigate Off-AIC2 that conducts auction-based matching and pricing between ESs and AgPs under predicted supply–demand conditions. Also, to ensure the effectiveness with explainability of our methodology, we offer theoretical analysis on individual rationality, incentive compatibility, and budget feasibility proofs under mild conditions.
  • Then, we turn to an online stage, wherein the scheduling process is modeled as a potential game among HuUs, MaUs, and heterogeneous ES/UAV, and develop a method called PG-BR, with convergence proof to a pure-strategy NE over controlled congestion and guaranteed fairness.
  • Extensive experiments based on real-world datasets demonstrate that our FUSION significantly reduces latency, improves resource utilization, and enhances robustness compared with state-of-the-art benchmarks.
II Literature Review
@@ II-A Service Provision in UAV-Assisted Edge Networks
UAV-assisted edge computing [8, 6, 7, 9] extends computation beyond static infrastructures, with resource scheduling coordinating communication, computation, and mobility. Existing studies primarily focus on efficiency, communication performance, and energy reduction in UAV-ground interactions. For instance, Wang et al. [6] designed a multi-agent deep reinforcement learning (MADRL)-based distributed scheduling scheme for intelligent reflecting surface (IRS)- and UAV-assisted wireless powered edge networks. Xu et al. [7] studied UAV-assisted internet of things (IoT) inference via DRL with an embedded resource allocation subroutine to minimize average inference delay. Li et al. [8] investigated federated edge intelligence for UAV-enabled IoT networks to jointly minimize UAV uploading energy and improve edge resource utilization. Zeng et al. [9] developed edge-server-assisted drone navigation via a DRL-based neural scheduler and network-aware containerized resource allocation to maximize a quality-of-navigation metric.
Although the above studies offer valuable insights, they predominantly rely on static or quasi-static geometric models that implicitly assume the UAV is always “nearby” and can provide immediate service. Consequently, they overlook critical factors such as UAV flight time, positional evolution, task arrival time, and the dynamic accessibility of ESs. Nevertheless, in practical arrival time, service sequencing, trajectory-dependent accessibility, and spatio-temporally varying workloads of UAVs play essential roles. These characteristics call for a transition from static coverage-based optimization toward spatio-temporal-aware frameworks.
II-B Auction-Driven Service Delivery over Distributed Edge Networks.
Auctions have been extensively investigated as an incentive-compatible paradigm over edge networks, enabling strategic users and heterogeneous service providers to transact multi-type services. For instance, Ren et al. [15] investigated truthful auctions for dependent task scheduling to maximize system utility in dynamic environments. Xu et al. [16] designed combinatorial auction–based distributed deep neural network (DNN) inference for heterogeneous edge–cloud networks to maximize social welfare. Li et al. [17] developed an online auction for artificial intelligence generated content (AIGC) task scheduling to optimize decisions on instances and time schedules. Lu et al. [18] presented auction-based cluster federated learning to improve convergence and balance energy consumption in edge networks.
While the above studies provide insights into incentive-compatible service provision in edge systems, they are mostly single-stage, reactive, and assume fixed resources and stable demand. This leads to high signaling overhead, latency, and energy costs, especially for large numbers of battery-constrained devices. In UAV-assisted edge computing, dynamic resources, bursty task arrivals, and strict delay/energy constraints call for time-efficient, spatio-temporal aware auction frameworks that reduce online interactions while handling real-world uncertainties.
@@
Our Work $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
II-C Game-empowered Service Provision over Distributed Edge Networks
Beyond auctions, potential-game-based approaches also explored distributed optimization strategies for edge service provisioning[12, 13]. Ding et al. [12] investigated end–edge–cloud computing architectures, formulating task assignment among different parties as potential games. Zhou et al. [13] modeled a potential game with latency–energy trade-off utilities to a pure-strategy NE. Hsieh et al. [19] proposed a knapsack potential game for task scheduling, aiming to balance QoS guarantees for different traffic priorities and overall cost-effectiveness. Wu et al. [20] developed a potential-game-based task scheduling scheme for multi-cell edge–cloud collaboration to a NE while minimizing users’ joint delay–energy cost.
Overall, potential-game-based studies show that carefully designed utilities and potential functions can achieve convergent, distributed scheduling equilibrium and reduce system costs. However, they generally consider a single, homogeneous user type, ignoring the distinct delay sensitivities and utility structures of HuUs and MaUs, as well as their interactions with heterogeneous ESs. Consequently, UAV-assisted platforms supporting human-machine coexistence, spatio-temporal task windows, and multi-party potential games remain largely unexplored. To highlight our contributions, Table I compares our work with representative existing studies.
III Overview and Modeling
III-A Overview of Our Methodology
We consider an air–ground integrated edge network, where all participants can be naturally grouped into two key categories: service demanders and service providers.
1) Service demanders (SDs). This includes users that generate tasks such as model training/inference [^2] We assume that each SD is associated with one task that requires computing services from SPs in every future trading round.. To capture fundamentally different QoS needs over human-machine coexistence, we further classify them into two groups:
  • HuUs, denoted by $\bm{\mathcal{U}}^{\mathrm{H}}=\{u_{1}^{\mathrm{H}},\dots,u_{|\bm{\mathcal{U}}^{\mathrm{H}}|}^{\mathrm{H}}\}$. HuUs are strongly sensitive to instantaneous service latency, as applications such as AR/VR, interactive gaming, real-time video conferencing, and intelligent driving can directly influence their experience and further emotion. Even slight increases in latency or service degradation may leave a heavy impact on their perceived utility.
  • MaUs, denoted by $\bm{\mathcal{U}}^{\mathrm{A}}=\{u_{1}^{\mathrm{A}},\dots,u_{|\bm{\mathcal{U}}^{\mathrm{A}}|}^{\mathrm{A}}\}$. MaUs correspond to autonomous devices such as robots, mapping vehicles, and sensing units. They primarily care about whether missions can be completed before a certain deadline, and are therefore less impacted by instantaneous latency, as long as overall task feasibility is preserved.
The coexistence of motion-sensitive HuUs and mission-oriented MaUs brings heterogeneous utility structures, imposing substantial challenges for optimized service delivery.
2) Service providers (SPs). This consists of two types of computing infrastructures:
@@ III-A Overview of Our Methodology
  • ESs, denoted by $\bm{\mathcal{S}}=\{s_{1},\dots,s_{|\bm{\mathcal{S}}|}\}$, which are geographically distributed across the map and provide low-latency services to nearby users. Each ES operates with a limited pool of orthogonal subcarriers and finite capacity, inherently constraining the number of users that can be concurrently served. During peak-demand periods, hotspot ESs are particularly susceptible to resource saturation, which can trigger substantial queuing delays or even outright admission blocking.
@@
  • AgPs, denoted by $\bm{\mathcal{V}}=\{v_{1},\dots,v_{|\bm{\mathcal{V}}|}\}$, with each formed by a vehicle with multiple UAVs charging on-board. Due to the limited power of UAVs, vehicles are firstly hired to traverse pre-planned routes, act as mobile carriers that transport UAVs to designated service locations. Upon approaching their contracted ESs, UAVs are deployed to offer local computing enhancing assistance. However, this paradigm also incurs non-trivial operational costs, including vehicle repositioning, UAV dispatch and launch procedures, flight energy consumption, and maintenance overhead.
Under this two-sided demand–supply structure, SDs compete for low-latency, reliable computing services, while SPs collaboratively offer computing resources under capacity and operational limitations. Nevertheless, in such a dynamic ecosystem, each ES can face uncertain service demand: the temporal and spatial distribution of service request peaks at each ES is highly stochastic, and the composition ratio between HuUs and MaUs in the service requests also varies dynamically over time.
Figure 1: Schematic of our FUSION over distributed air–ground integrated edge network.
Against this background, we develop hybrid and coupled stages. As illustrated in Fig. 1, the offline stage first trains an LNN to perform multi-step spatio-temporal forecasting of future service demand at each ES. Guided by this, we jointly plan vehicular visiting routes and the initial deployment of UAVs, while simultaneously designing auction-driven incentive for trading contracts between ESs and AgPs. In online Stage, interactions among SDs (e.g., HuUs and MaUs), and heterogeneous SPs (e.g., ESs and UAVs) are modeled by forming a potential game. By constructing an ordinal potential function to capture the global utility, we allow each player updates its strategy locally and the overall performance converges to a pure-strategy NE, yielding a near-global optimal solution.
In summary, our ultimate goal is to design a sustainable, intelligent, and incentive-compatible framework for distributed edge service provision, that organically integrates forecasting-driven offline contracting with game-based online scheduling, so as to balance service performance and operational cost, further achieving multi-party win–win over human–machine coexistence.
III-B Basic Modeling
In the following, we offer detailed modelings for players.

### v2

@@ Abstract
In this paper, we introduce a first-of-its-kind forecasting-driven, incentive-inherent service provisioning framework for distributed air-ground integrated networks that explicitly accounts for human-machine coexistence. In our framework, vehicular-UAV agent pairs (APs) are proactively dispatched to overloaded hotspots to augment the computing capacity of edge servers (ESs), which in turn gives rise to a set of challenges that we jointly address: highly uncertain spatio-temporal workloads, spatio-temporal coupling between road traffic and UAV capacity, forecast-driven contracting risks, and heterogeneous quality-of-service (QoS) requirements of human users (HUs) and machine users (MUs). To address these challenges, we propose FUSION, a two-stage optimization framework, consisting of an offline stage and an online stage. In the offline stage, a liquid neural network-powered module performs multi-step spatio-temporal demand forecasting at distributed ESs, whose outputs are exploited by an enhanced ant colony optimization-based routing scheme and an auction-based incentive-compatible contracting mechanism, to jointly determine ES-AP contracts and pre-planned service routes. In the online stage, we formulate the congestion-aware task scheduling as a potential game among HUs, MUs, and heterogeneous ES/UAVs, and devise a potential-guided best-response dynamics algorithm that provably converges to a pure-strategy Nash equilibrium. Experiments on both synthetic and real-world datasets show that FUSION consistently achieves higher social welfare and improved resource utilization, while maintaining latency and energy costs comparable to state-of-the-art baselines and preserving individual rationality, budget balance, and near-truthfulness.
@@
Distributed edge computing, liquid neural network, auction, potential games, heterogeneous network users.
@@ I Introduction
Recent years have witnessed a new generation of intelligent applications such as real-time distributed video analytics, intelligent transportation systems, and large-scale mobile crowdsensing, flourishing across smart cities [1, 2]. Although edge servers (ESs) can provide computing capabilities for these intelligent applications at the network edge, their limited local resources make them susceptible to resource shortages and task congestion during peak demand periods [3].
To address the limitations of ESs, low-altitude edge computing enabled by unmanned aerial vehicles (UAVs) has received significant attention [4, 5]. Unlike static ESs, UAVs provide high mobility and flexibility, allowing them to be rapidly dispatched to hotspot regions to provide the needed computing resources. In this work, we highlight an important yet often-overlooked aspect of low-altitude edge systems: most existing studies implicitly assume that UAVs can always fly directly to target regions[8, 7, 9, 6], without considering their deployment origins. This is despite the fact that long travel distances due to poorly planned deployment origins can easily deplete UAVs’ batteries before reaching target locations. To fill this gap, we propose an air-ground cooperative edge computing paradigm: during periods of intensive demand, a set of ground vehicles are used to transport UAVs (vehicles and UAVs are collectively referred to “agent pairs (APs)”) to the vicinities of overloaded ESs. UAVs are then launched from strategically selected locations to deliver temporary computational augmentation for ESs. Despite its promising potential, operationalizing this hybrid AP-ES deployment paradigm introduces several challenges, which must be addressed:
(A) Spatio-temporal demand uncertainty. Edge computing demands exhibit strong spatio-temporal variability, and the coexistence of human users (HUs) and machine users (MUs) further complicates the management of distributed workloads. This variability makes it difficult to determine when, where, and how UAV resources should be dispatched. In such networks, static provisioning or purely reactive scheduling often leads to mismatches between supply and demand, resulting in congestion at emerging hotspots and resource underutilization in low-demand regions [13, 12]. The key challenge therefore lies in accurately forecasting spatio-temporal demand patterns from historical requests and contextual information, so as to enable proactive UAV deployment.
(B) Spatio-temporal coupling of vehicular mobility and UAV deployment. Supporting geo-distributed ESs requires APs to rely on vehicular mobility to reach target locations, which is constrained by road topology and traffic conditions. Moreover, the number of UAVs that can be deployed by a vehicle is tightly coupled with its visiting sequence and dwell time at each ES, since UAV dispatching, service execution, and recovery jointly consume time and onboard resources. Consequently, vehicle–UAV joint deployment becomes a highly coupled optimization problem that must simultaneously account for travel costs, service time windows, UAV dispatching logistics, and spatially distributed service rewards.
(C) Predictive contracting and incentive-compatible trading. In the ES–AP hybrid network, APs operated by private entities must be compensated by ES operators through advance contracts that commit service capacity and pricing for future peak-demand periods before demand realization. However, forecasting errors and dynamic supply–demand variations introduce significant risks for both parties: ESs may suffer unmet demand, while APs may incur revenue loss due to resource misallocation. The central challenge is thus to design a reliable offline trading mechanism that is individually rational, incentive compatible, and budget feasible, while operating prior to demand realization and service execution [14].
(D) Online network coordination under human–machine coexistence. HUs and MUs compete for shared resources provided by local ESs and leased UAVs, yet exhibit fundamentally different QoS requirements. HUs are typically delay-sensitive due to interactive applications, whereas MUs emphasize timely task completion before deadlines. The resulting challenge is to design fair and efficient distributed resource scheduling mechanisms that jointly serve HUs and MUs while ensuring stable operation and provable convergence guarantees.
@@
The above four challenges form the motivation of our work. Specifically, to address them, we develop a forecasting-driven agent scheduling and incentive optimization (FUSION) framework for distributed service provisioning. In FUSION, to address challenge (A), we design a proactive spatio-temporal demand prediction via liquid neural networks (Pro-LNN), which transforms historical service traces and contextual features into multi-step workload forecasts. Then, to address challenge (B), we propose an enhanced ant colony optimization for vehicular route planning (eACO-VRP) based on the predicted demand. Afterwards, to address challenge (C), we introduce offline auction-based incentive-compatible contracting (Off-AIC2) that matches ESs with proper APs and obtains payments and rewards. Finally, to address challenge (D), we formulate an efficient online potential game among HUs, MUs, and ES/UAV, and propose the potential game-based best-response dynamics (PG-BRD) method, which integrates potential-guided dual decomposition with asynchronous best-response dynamics to drive the system towards a pure-strategy Nash equilibrium (NE) under congestion and fairness constraints. To sum up, our major contributions can be summarized as follows:
  • We present one of the first air-ground integrated edge network models that jointly captures vehicular-UAV APs, ESs, and heterogeneous human-machine (i.e., HU-MU) computing demands within a single framework.
  • We formulate ES–AP–user service provisioning by accounting for spatio-temporal mobility constraints, stochastic workload evolution, and strategic agent behaviors. We unveil that the offline ES–AP contract design admits a combinatorial mixed-integer formulation, while the online interactions between service demanders (HUs and MUs) and providers (ESs and APs) leads to a finite ordinal potential game with a combinatorial action space. To tackle these formulations, we develop a two-stage solution consisting of a forecast-driven offline stage and a game-theoretic online stage.
  • For the offline stage, we develop a multi-step spatio-temporal demand forecasting framework via Pro-LNN with executable vehicle-UAV route planning through an enhanced ACO-based algorithm (i.e., eACO-VRP). To our knowledge, this represents one of the first applications of the emerging LNN paradigm to spatio-temporal workload forecasting for proactive decision-making in air-ground integrated networks.
  • To address forecast-driven risk and strategic interactions in the offline stage, we propose Off-AIC2, a novel auction-based offline ES-AP contracting mechanism that incorporates demand forecasts into demand-supply matching and resource pricing, with provable guarantees of individual rationality, incentive compatibility, and budget feasibility.
  • For the online stage, we model real-time task scheduling as a multi-player potential game involving HUs, MUs, and heterogeneous ES/UAV resources. We then design PG-BRD, a potential-guided best-response algorithm with convergence guarantees to a pure-strategy NE.
  • Through extensive evaluations, we show that FUSION consistently delivers higher social welfare while keeping the delay and energy costs on par with existing baselines.
II Related Work
@@ II-A Service Provision in UAV-Assisted Edge Networks
There exist an extensive literature on improving the communication/computation performance and reducing the energy footprint of UAV-assisted edge computing systems [8, 6, 7, 9]. For instance, Wang et al. [6] designed a multi-agent deep reinforcement learning (DRL)-based resource scheduling scheme for UAV-assisted wireless edge networks; Xu et al. [7] studied minimizing the average inference delay over UAV-assisted computing systems via DRL-based resource allocation; Li et al. [8] investigated distributed/federated edge intelligence for UAV-assisted networks, while optimizing the UAV energy consumption and resource utilization; Zeng et al. [9] introduced an ES-assisted UAV navigation system that integrates DRL-based decision making with network-aware containerized resource allocation to maximize the quality of navigation.
Although the above studies offer valuable insights, they predominantly presume that UAVs are always nearby and can provide immediate service, often without explicitly modeling their underlying deployment units or origins. As a result, the joint treatment of the four challenges summarized in Sec. I remains largely unexplored, an effort undertaken in this work.
II-B Auction-Driven and Game-Theoretic Service Provisioning
Auctions have been widely used for resource provisioning from service providers to users. For instance, Ren et al. [15] investigated truthful auctions for dependent task scheduling over dynamic edge networks; Xu et al. [16] designed combinatorial auction-based deep neural network (DNN) inference for heterogeneous edge-cloud networks to maximize social welfare; Li et al. [17] developed an online auction for generative content scheduling; Lu et al. [18] proposed auction-based distributed/federated learning to improve the distributed model training speed and energy consumption in edge networks. While the above studies provide valuable insights, they are mostly reactive (i.e., decisions are made after demand realization), and typically assume fixed resources and stable demand. In contrast, UAV-assisted edge computing involves dynamic resources and stringent delay and energy constraints, calling for spatio-temporally-aware auctions with minimal online interactions between service providers and end users while accounting for uncertainties in both resource availability and demand. This gap is addressed by our Off-AIC2 framework.
Beyond auctions, potential games have also been explored as distributed optimization strategies for service provisioning[12, 13]. For example, Ding et al. [12] investigated task processing across user–edge–cloud networks as potential games; Zhou et al. [13] modeled computation offloading as a potential game with latency–energy trade-offs; Hsieh et al. [19] proposed a knapsack potential game for task scheduling; Wu et al. [20] studied task scheduling for multi-cell edge–cloud networks through potential games with NE guarantees. Despite their notable contributions, the above works typically consider a single/homogeneous user type, overlooking the fundamentally different QoS requirements of HUs and MUs, as well as their interactions with heterogeneous ESs. Consequently, the effectiveness of potential games in UAV-assisted platforms that support human-machine coexistence, spatio-temporal task windows, and multi-party interactions remains largely unexplored, an effort addressed via our developed PG-BRD.
To further highlight our contributions, Table I compares our approach with representative existing studies.
III Overview and Modeling
III-A Overview of Our Methodology
We consider an air-ground integrated network, wherein nodes are partitioned into service demanders and service providers.
1) Service demanders (SDs): SDs are users that generate tasks such as model training/inference [^1] We assume that each SD is associated with one task that requires computing services from SPs in every future trading round.. To capture different QoS needs under human-machine coexistence, we further classify SDs into two categories:
  • HUs, denoted by $\bm{\mathcal{U}}^{\mathrm{H}}=\{u_{1}^{\mathrm{H}},\dots,u_{|\bm{\mathcal{U}}^{\mathrm{H}}|}^{\mathrm{H}}\}$ are sensitive to instantaneous service latency, as applications such as AR/VR, interactive gaming, real-time video conferencing, and intelligent driving are interactive and experience-driven, where even a slight change in latency or QoS can leave a pronounced impact on their perceived utility.
  • MUs, denoted by $\bm{\mathcal{U}}^{\mathrm{M}}=\{u_{1}^{\mathrm{M}},\dots,u_{|\bm{\mathcal{U}}^{\mathrm{M}}|}^{\mathrm{M}}\}$, correspond to autonomous devices such as robots, mapping vehicles, and sensing units. MUs are primarily concerned about whether their computation tasks can be completed before a certain deadline, and are therefore less impacted by instantaneous latency, as long as their overall task feasibility is preserved (e.g., a robot’s image-analysis task is considered feasible as long as the inference results are obtained before the deadline).
@@ III-A Overview of Our Methodology
This Work $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
2) Service providers (SPs): We consider two types of SPs:
@@
  • ESs, denoted by $\bm{\mathcal{S}}=\{s_{1},\dots,s_{|\bm{\mathcal{S}}|}\}$, are geo-distributed across the map and provide low-latency services to nearby users. Each ES operates with a limited pool of orthogonal subcarriers and finite resource capacity, inherently constraining the number of users that can be concurrently served. During peak demand periods, hotspot ESs are particularly susceptible to resource saturation, which can trigger substantial queuing delays or even admission blocking.
@@
  • APs, denoted by $\bm{\mathcal{V}}=\{v_{1},\dots,v_{|\bm{\mathcal{V}}|}\}$, are formed by vehicles carrying multiple on-board charging UAVs. Due to the limited power of UAVs, vehicles first traverse pre-planned routes to transport UAVs to designated service/ES locations. Upon arriving at their contracted ESs, UAVs are deployed to offer extra computing resources. This procedure incurs operational costs, including vehicle repositioning, UAV dispatching/launching procedures, flight energy consumption, and maintenance overhead.
Under this two-sided demand-supply structure, SDs compete for reliable computing services, while SPs collaboratively offer computing resources under their capacity and operational limitations. Nevertheless, in such a dynamic and distributed ecosystem, ESs face uncertain service demand: the temporal and spatial distribution of service request peaks across ESs is highly stochastic, and the composition ratio between HUs and MUs in the service requests also varies over time. To capture the time-varying behavior, a practical transaction is discretized into $T$ time-slots, with index $t$, i.e., $t\in\left\{0,1,2,...,T\right\}$. Further, to enable resource provisioning between SDs and SPs, we adopt a two-stage resource allocation framework that explicitly separates an offline stage from an online stage. From the ES perspective, the offline stage operates before the actual trading round, when only historical records and statistical demand patterns are available. At this stage (detailed in Sec. IV), the identities and tasks of HUs and MUs in the next trading round are unknown, making it impossible to pre-plan precise AP-ES matchings. Accordingly, in the offline stage of Fig. 1, we train Pro-LNN to forecast future service demand for each ES, and, based on these predictions, jointly design AP-to-ES visiting routes, UAV deployment plans, and auction-based contracts between ESs and APs (i.e., Off-AIC2).
Figure 1: Schematic of FUSION over an air–ground integrated network.
In contrast, the online stage starts (detailed in Sec. V) when real computing tasks arrive and the actual identities and locations of SDs become known. With full knowledge of active participants and their instantaneous task requirements, the system performs real-time task–resource scheduling. In this stage, we model the interactions among heterogeneous SPs and SDs as a potential game and construct an ordinal potential that captures the social welfare. Each SD then updates its task offloading decision and each SP adjusts its service admission in a fully distributed manner, where the resulting potential-guided best-response dynamics converge to a pure-strategy NE, yielding an online and congestion-aware provisioning of resources.
III-B Modeling of SDs, SPs, and Contracts
Hereafter, we provide detailed modeling of SPs and SDs.